# Group-On: Boosting One-Shot Segmentation with Supportive Query

### v1

@@ Abstract
One-shot semantic segmentation aims to segment query images given only ONE annotated support image of the same class. This task is challenging because target objects in the support and query images can be largely different in appearance and pose (i.e., intra-class variation). Prior works suggested that incorporating more annotated support images in few-shot settings boosts performances but increases costs due to additional manual labeling. In this paper, we propose a novel approach for ONE-shot semantic segmentation, called Group-On, which packs multiple query images in batches for the benefit of mutual knowledge support within the same category. Specifically, after coarse segmentation masks of the batch of queries are predicted, query-mask pairs act as pseudo support data to enhance mask predictions mutually, under the guidance of a simple Group-On Voting module. Comprehensive experiments on three standard benchmarks show that, in the ONE-shot setting, our Group-On approach significantly outperforms previous works by considerable margins. For example, on the COCO-20i dataset, we increase mIoU scores by 8.21% and 7.46% on ASNet and HSNet baselines, respectively. With only one support image, Group-On can be even competitive with the counterparts using 5 annotated support images.
@@
Figure 1: (a) To deal with queries from user, the common 1-shot setting uses one support image annotated by one annotator. (b) The common few-shot setting uses a few support images annotated by many extra annotators. (c) Our one-shot Group-On handles queries in bulk based on ONE labeled support without increasing extra annotators. All the queries are interchangeable and achieve mutual complementary benefits. Refer to caption and Refer to caption represent the input of support and pseudo support (i.e. supportive queries & pseudo masks), repectively. The bar charts on the bottom show the average segmentation performances on PASCAL-5i and COCO-20i datasets. Obviously, our one-shot Group-On even reaches a competitive level with the counterparts under the 5-shot setting.
1 Introduction
@@
Recently, few-shot learning (FSL) has garnered significant attention due to its remarkable success in reducing the dependence on data annotations during the development of deep learning (DL) models. This paradigm offers an efficient way to deploy a DL model that can generalize well to new and unseen data by using only a sparing amount of category-identical reference data with manual labels. This is especially advantageous for tasks that require costly manual annotations, such as pixel-wise segmentation of images. In a few-shot segmentation (FSS) practice, users require the segmentation model to segment unseen images (i.e., query) with guidance from one or several references (i.e., support) of the same class. Apparently, “one-shot setting” is the most annotation-effective solution, requiring the resource-intensive pixel-wise labeling of just one image. However, constrained by the limited information derived from just one manual annotation and the intrinsic diversity of the objects to be segmented, its performance often falls behind in other few-shot scenarios [1].
@@
Inspired by this, we try to integrate the merits of both one-shot and few-shot settings for fewer annotations and higher performances. Interestingly, a large quantity of query images from the same category can also offer intra-class multiplicity intuitively and thus provide mutual complementary support and benefits, which, however, has been largely neglected in previous research. In this paper, we focus on one-shot segmentation that fixes the number of support images to only ONE. We propose a novel approach for one-shot segmentation, called Group-On, which simultaneously handles multiple queries input by users without increasing extra annotation costs and obtains better performance via mutual assistance among the queries, as seen in Fig. 1(c). For better understanding of this approach, we define the query selected to be predicted as the host query, while the rest queries as supportive queries. Supportive queries and their intermediate coarse masks, predicted with the only support image, serve as pseudo support data to facilitate segmentation of the host query. In other words, the host query is segmented by the supports from one real support image and multiple pseudo ones (i.e., supportive queries), resulting in multiple coarse prediction candidates. Eventually, such candidate masks for the host query are integrated together to yield the final prediction.
@@
Note that the host and supportive queries hold an equal and interchangeable role. Each one among a group of query images can be the host query to be segmented finally, and all the query images offer favorable information to each other simultaneously. Due to the need for testing many query samples after the deployment of a model in real-world scenarios, it is reasonable and feasible to collect and use various queries and apply Group-On approach for reciprocal help. Since this approach can be viewed as somewhat similar to a “group buying” activity where the individuals benefit from the collective effort, we refer to it as “Group-On”. Experiments show that our approach in the 1-shot setting is even competitive with methods in the 5-shot setting which requires much more time-consuming manual annotations.
In summary, our main contributions are as follows:
@@
  • We rethink the real-world application scenarios and propose Group-On, a simple, novel, and effective training approach for one-shot segmentation. Group-On simultaneously processes category-identical queries uploaded by users to leverage the rich object pattern variety to address the intra-class variation issue.
  • We propose a Group-On Voting module to assign voting weights to candidate masks based on attention correlations between host and supportive query pairs, which promotes steady training as well. Furthermore, in the case when not many testing samples are available, we replace attention voting weights with arithmetic mean for lower computational burden.
@@
  • We perform extensive experiments and analysis on three challenging FSS benchmarks: PASCAL-5i [5], COCO-20i [6], and FSS-1000 [7]. The results show that the performances of one-shot segmentation baselines with our Group-On are comparable to the counterparts of conventional 5-shot segmentation.
2 Related Work
It has been widely acknowledged that OSLSM [8] represents the pioneering work in one-shot segmentation. Inspired by similar architectures for few-shot classification, OSLSM is comprised of a condition branch and a segmentation branch. After that, numerous methods were put forward to improve the performance of FSS, most of which investigated interactions between support and query images while the others focused on details inside either the support or query data.
2.1 Interactions between Support and Queries
Many studies [4, 9, 10] sought to design high-capacity interaction modules between the support and query data. It was common to expand features in multiple scales and layers [11, 12] to enhance in-depth interplay. Based on the Transformer architecture, CyCTR [13], DCAMA [14], and HDMNet [15] developed cross-attention intercommunication while some others [16, 3, 17, 18, 19] did not. Besides, there were some attempts to generate high-dimensional feature correlation tensors, such as HSNet [20], ASNet [1], VAT [21], and DACM [22].
2.2 Support Exploration
Due to the scarcity of reference information, some studies took advantage of the support data in various perspectives (e.g., background and boundaries). PANet [23] proposed to take the predicted query data as pseudo references to segment support images in turn, which aligns both support and query clues and guides each other. As the knowledge of objects from base classes in the background can be ample, it was investigated [24, 25, 26] to use such knowledge to assist novel class predictions. HM [27] and MSI [28] employed two masking techniques to segment more accurate target object boundaries. In addition, IPRNet [29] succeeded in reducing inter-class similarity, but demanded more reference labels even in training under the 5-shot setting, while most methods used five support annotations just in testing. Recently, MIANet [30] introduced supplementary category labels to leverage the general semantic information in a multi-modal manner.
2.3 Query Exploration
For query exploration, DPNet [31] added pseudo-prototypes extracted from the query foreground to the original support prototypes for refinement. SSP [32] suggested that the locality of query masks, which were predicted in high confidence, could aid the support data to segment the leftover low-confidence areas. Unlike the supervised training mode mentioned above, QGNet [33] trained an extractor to approximately locate the targets via global-local contrastive learning.
The known methods on query exploration mainly aimed to make full use of what is inside a single query image to guide the segmentation. In contrast, our Group-On approach exploits directly increasing the query amount to enrich semantic knowledge in an intuitive way. Group-On is more capable and simpler, and can be incorporated seamlessly with plentiful FSS baselines.
@@ 2.3 Query Exploration
Figure 2: Illustrating Group-On in the one-shot setting. The left is an overview of the proposed approach and we present just one supportive query here for better understanding. Different from the conventional few-shot setting, a group of query images along with one support image is input into the model. After the coarse masks of all the query images are segmented, the supportive image-mask pairs function as pseudo support to segment the host query again. Finally, the Group-On Voting module assigns voting weights to the segmentation masks of the host query and produces the final prediction result. Note that the host and supportive queries play an equal and interchangeable role. For ease of understanding, we present the probability maps in the form of segmentation masks, while they are actually two-dimensional floating-point matrices. The top right corner is details of Group-On Voting module. Note that the input $\textit{{F}}^{\text{supp}}_{1}$, $\textit{{F}}^{\text{supp}}_{2}$, and $\textit{{F}}^{\text{host}}$ are extracted from a middle layer of the feature extractor.
3 Methodology
3.1 Problem Setting
Few-shot segmentation (FSS) aims to segment images with limited annotated examples. To improve generalization [34], we adopt the prevalent meta-training paradigm called episodic training [35], following previous work [20, 21]. Concretely, a dataset is divided into two class-disjoint sets $\mathcal{D}_{train}$ and $\mathcal{D}_{test}$, each of which contains many episodes (an episode usually consists of a query and a support set). What is very particular in FSS is that in an episode, there exist a support set $\mathcal{S}=\{(\textbf{{I}}^{s},\textbf{{M}}^{s})\}$ and a query set $\mathcal{Q}=\{(\textbf{{I}}^{q},\textbf{{M}}^{q})\}$ from the same class, where $\textbf{{I}}^{z}$ and $\textbf{{M}}^{z}$ represent raw images and their binary masks.
@@ 3.1 Problem Setting
In fact, each episode instantiates an $N$-way $K$-shot segmentation task. That is, we select $K$ support image-label pairs per category from $N$ different categories to construct $\mathcal{S}$. Current approaches typically utilize one support image-label pair for training whereas one (i.e., 1-way 1-shot) or five (i.e., 1-way 5-shot) for testing, and one query image during both stages. The number of support images matters since it determines the burden of labeling. We always seek to obtain similar performances via as few support images as possible and preferably only one. That is why many models [8, 36] freeze the number of support images to one for training.
@@
In this work, apart from the 1-way setting following [20], we propose and undertake a 1-support $C$-query setting ($C>1$) extended from the conventional settings. That is, the amount of support images is ONE but we use multiple query images to support each other in the process.
3.2 Overview of Group-On
The overall pipeline of Group-On is presented in Fig. 2. Basically, it performs one-shot segmentation in three main steps. First, after inputting one support image and query images in a group, we obtain corresponding $C$ predicted masks based on the support data, among which one acts as a candidate mask while the others function as pseudo support masks. Second, given $C-1$ pseudo support image-mask pairs, the parameter-sharing segmentation head predicts the remaining $C-1$ candidate segmentation masks. Finally, the Group-On Voting module receives and integrates $C$ candidate masks to produce the final prediction. Below we first describe our approach under the 1-support $C$-query setting, and then detail the Group-On Voting module.
3.3 Group-On Process under One-Shot Setting
@@
Aiming to maximize the semantic information of different patterns in the same class as much as possible, we use a group of $C$ query images and one support image-label pair as input to the model. The pivotal competitive advantage of Group-On is in utilizing the group of query images for mutual knowledge complement. Among the $C$ query images in the group, the final predicted image is called the host query while the rest images are supportive queries. For simple illustration, in this section we assume that $C=2$, and the host query and its feature are denoted as $\textit{{I}}^{q}_{1}$ and $\textit{{F}}^{q}_{1}$, while the supportive ones as $\textit{{I}}^{q}_{2}$ and $\textit{{F}}^{q}_{2}$.
@@
The entire data stream goes through the segmentation head $\Psi$ twice. In the first time, the support feature-mask pair ($\textit{{F}}^{s}$, $\textit{{M}}^{s}$) and the query features $\textit{{F}}^{q}_{i}$’s are input into the model to generate the initial coarse masks, as:
$\hat{\textit{{M}}}^{q}(i)=\Psi(\textit{{F}}^{s},~{}\textit{{M}}^{s},~{}\textit{{F}}^{q}_{i}),$ (1)
@@
where $\hat{\textit{{M}}}^{q}(i)$ denotes the coarse mask predicted using the real support data ($\textit{{F}}^{s}$, $\textit{{M}}^{s}$), $i=1,2,\ldots,C$ ($C=2$ here).
Among the initial coarse masks, $\hat{\textit{{M}}}^{q}(1)$ is used as a candidate mask (for the host query), that is:
$\bar{\textit{{M}}}^{q}(1,1)=\hat{\textit{{M}}}^{q}(1).$ (2)
@@
$\hat{\textit{{M}}}^{q}(2)$, $\ldots$,… , $\hat{\textit{{M}}}^{q}(C)$ and their corresponding supportive queries are used as pseudo support data, which are input into the segmentation head again to compute the other candidate masks, as:
$\bar{\textit{{M}}}^{q}(1,j)=\Psi(\textit{{F}}^{q}_{j},~{}\hat{\textit{{M}}}^{q}(j),~{}\textit{{F}}^{q}_{1}),j=2,\ldots,C,$ (3)
@@
where $\bar{\textit{{M}}}^{q}(1,j)$ is the candidate mask of $\textit{{I}}^{q}_{1}$ derived from the supportive feature-mask pair ($\textit{{F}}^{q}_{j},~{}\hat{\textit{{M}}}^{q}(j)$), $j=2,\ldots,C$. Hence, we obtain $C$ candidate masks $\bar{\textit{{M}}}^{q}(1,1),\ldots,\bar{\textit{{M}}}^{q}(1,C)$, produced from true and pseudo references in distinct appearances but from the same class, thus increasing the object diversity. The final result ${\textit{{M}}}^{q}_{1}$ is computed as:
${\textit{{M}}}^{q}_{1}=\Upsilon(\bar{\textit{{M}}}^{q}(1,1),\ldots,\bar{\textit{{M}}}^{q}(1,C)),$ (4)
where $\Upsilon$ is the Group-On Voting module to be presented in the next subsection.
@@
It is noteworthy that under the 1-shot setting, all the query images are treated equivalently when applying Group-On. That is, the host query is not limited only to $\textit{{I}}^{q}_{1}$ but must be every query image, and the remaining supportive queries are processed in the same way as what is described above. This process is able to attain mutual knowledge complement. The reason for this is that it allows to make the most out of clues hidden in the query data and promote efficiency.
3.4 Group-On Voting Module
To maximize the utilization of the candidate masks, we develop the Group-On Voting module to vote on each candidate based on the correlations between the host query and the supportive queries/support image for advanced mutual information promotion, which is a process of assigning voting weights. The support image functions similarly as the supportive queries in the Group-On Voting module. Thus, the supportive queries and support image are collectively referred to as supportive query data below for simplicity.
For meta-training, the quality of the coarse masks $\hat{\textit{{M}}}^{q}(i)$, $i=1,\ldots,C$, is woefully poor at the beginning due to non-convergence, which demands stricter filtering on the candidates predicted with the coarse masks. As shown in the top right corner of Fig. 2, we leverage the cross-attention correlations between the host-supportive query pairs to determine how much a candidate mask contributes to the final result during integration. First, from a middle layer of the feature extractor (ResNet50 [37] or ResNet101 [37]), we obtain feature maps $\textit{{F}}^{\text{host}}_{1},\textit{{F}}^{\text{supp}}_{i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ of the host and supportive queries, for each $i=1,\ldots,C$. Second, we squeeze the feature maps through their spatial dimensions $H^{\prime}\times W^{\prime}$ with global average pooling to produce feature vectors $\textit{{v}}^{\text{host}}_{1},\textit{{v}}^{\text{supp}}_{i}\in\mathbb{R}^{1\times 1\times C^{\prime}}$, $i=1,\ldots,C$. Then following the general form of attention [38], the voting weight $w_{i}$ is calculated as:
$w_{i}=\frac{\textit{{v}}^{\text{host}}_{1}\otimes(\textit{{v}}^{\text{supp}}_{i})^{\mathsf{T}}}{\sqrt{d}},\ \ i=1,\ldots,C,$ (5)
where $\otimes$ is the cross product and $d$ is a hyper-parameter. Here we set $d=1$. Finally, the voting weights $W=\{w_{i}\}_{i=1}^{C}$ activated by the Softmax function are assigned to the $C$ candidate masks $\bar{\textit{{M}}}^{q}(1,i)$’s for the final prediction:
${\textit{{M}}}^{q}_{1}=\frac{1}{C}\sum_{i=1}^{C}\bar{\textit{{M}}}^{q}(1,i)*w_{i},\ \ i=1,\ldots,C,$ (6)
where ${\textit{{M}}}^{q}_{1}$ is the final prediction described in Eq. (4).
As for meta-testing, the model is able to segment relatively decent coarse masks after convergence. If we need to reduce the computational requirements of the filtering, we can turn our attention to attaining less computation and faster inference. Thus, the arithmetic mean is employed, as:
${\textit{{M}}}^{q}_{1}=\frac{1}{C}\sum_{i=1}^{C}\bar{\textit{{M}}}^{q}(1,i),\ \ i=1,\ldots,C,\vspace{-0.5em}$ (7)
We should emphasize that the host and supportive queries play equal and exchangeable roles, and we take $\textbf{{I}}^{q}_{1}$ as the host query just as an illustrating example here. Different from the few-shot setting, we train and test query images in groups by aiding each other concurrently in the one-shot setting. As the training process progresses, the performance of the model increases gradually and boosts the coarse mask quality, which eventually enhances the final prediction.
@@ 3.4 Group-On Voting Module
Table 1: Segmentation performances on the COCO-20i (left) and PASCAL-5i (right) datasets under the 1-support $C$-query setting. HSNet5 and ASNet5 represent the results of the baseline models HSNet and ASNet under the conventional 5-support setting, while the rest results are all in the 1-support setting. We highlight the best results in bold and underline the second best results for better viewing.
@@
5-support HSNet5 [20] 43.30 51.30 48.20 45.00 46.90 70.70 70.30 73.20 67.30 67.10 69.50 80.60
ASNet5 [1] 47.60 50.10 47.70 46.40 47.90 71.60 72.60 74.30 65.30 67.10 70.80 80.40
1-support SSP [32] 35.50 39.60 37.90 36.70 37.40 - 60.50 67.80 66.40 51.00 61.40 -
DCAMA [14] 41.90 45.10 44.40 41.70 43.30 69.50 67.50 72.30 59.60 59.00 64.60 75.70
VAT [21] 39.00 43.80 42.60 39.70 41.30 68.80 67.60 71.20 62.30 60.10 65.30 77.40
FECANet [4] 38.50 44.60 42.60 40.70 41.60 69.60 69.20 72.30 62.40 65.70 67.40 78.70
QCLNet [39] 39.80 45.70 42.50 41.20 42.30 - 65.20 70.30 60.80 61.00 64.30 -
MCE [40] 42.13 48.30 43.68 42.76 44.22 - 65.30 71.22 66.17 61.04 65.93 78.10
HSNet [20] 36.30 43.10 38.70 38.70 39.20 68.20 64.30 70.70 60.30 60.50 64.00 76.70
ASNet [1] 41.50 44.10 42.80 40.60 42.40 68.80 68.90 71.70 61.10 62.70 66.10 77.70
HSNet + Group-On 41.94 48.93 46.66 43.14 45.17 (5.97 $\uparrow$ ) 71.80 (3.60 $\uparrow$ ) 71.32 73.53 61.81 64.18 67.71 (3.71 $\uparrow$ ) 78.75 (2.05 $\uparrow$ )
ASNet + Group-On 46.34 49.59 48.60 46.73 47.82 (5.42 $\uparrow$ ) 72.60 (3.80 $\uparrow$ ) 73.07 73.87 61.73 67.71 69.10 (3.00 $\uparrow$ ) 80.05 (2.35 $\uparrow$ )

### v2

@@ Abstract
One-shot semantic segmentation aims to segment query images given only ONE annotated support image of the same class. This task is challenging because target objects in the support and query images can be largely different in appearance and pose (i.e., intra-class variation). Prior works suggested that incorporating more annotated support images in few-shot settings boosts performances but increases costs due to additional manual labeling. In this paper, we propose a novel and effective approach for ONE-shot semantic segmentation, called Group-On  [^1] https://github.com/diaoshaoyou/Group-On, which packs multiple query images in batches for the benefit of mutual knowledge support within the same category. Specifically, after coarse segmentation masks of the batch of queries are predicted, query-mask pairs act as pseudo support data to enhance mask predictions mutually. To effectively steer such process, we construct an innovative MoME module, where a flexible number of mask experts are guided by a scene-driven router and work together to make comprehensive decisions, fully promoting mutual benefits of queries. Comprehensive experiments on three standard benchmarks show that, in the ONE-shot setting, Group-On significantly outperforms previous works by considerable margins. With only one annotated support image, Group-On can be even competitive with the counterparts using 5 annotated images.
@@
Figure 1: (a) To deal with queries from user, the common 1-shot setting uses one support image annotated by one annotator. (b) The common few-shot setting uses a few support images annotated by many extra annotators. (c) Our one-shot Group-On handles queries in bulk based on ONE labeled support without increasing extra annotators. All the queries are interchangeable and achieve mutual complementary benefits. Refer to caption and Refer to caption represent the input of support and pseudo support, respectively. The bar charts of segmentation performances on the bottom show that our one-shot Group-On even reaches a competitive level with the counterparts in 5-shot setting.
I Introduction
@@
Recently, few-shot learning has garnered significant attention due to its remarkable success in reducing the dependence on data annotations during the development of deep learning (DL) models. This paradigm offers an efficient way to deploy a DL model that can generalize well to new and unseen data by using only a sparing amount of category-identical reference data with manual labels. This is especially advantageous for tasks that require costly manual annotations, such as pixel-wise segmentation of images. In a few-shot segmentation (FSS) practice, users require the segmentation model to segment unseen images (i.e., query) with guidance from one or several references (i.e., support) of the same class. Apparently, “one-shot setting” is the most annotation-effective solution, requiring the resource-intensive pixel-wise labeling of just one image. However, constrained by the limited information derived from just one manual annotation and the intrinsic diversity of the objects to be segmented, its performance often falls behind in other few-shot scenarios [1].
@@
Inspired by this, we try to integrate the merits of both one-shot and few-shot settings for fewer annotations and higher performances. Interestingly, a large quantity of query images from the same category can also offer intra-class multiplicity intuitively and thus provide mutual complementary support and benefits, which has been largely neglected in previous research yet. In this paper, we focus on one-shot segmentation that fixes the number of support images to only ONE. We propose a novel and effective approach for one-shot segmentation, called Group-On (Fig. 1(c)), which simultaneously handles multiple queries input by users without increasing extra annotation costs and obtains better performance via mutual assistance among the queries. For better understanding of Group-On, we define the query selected to be predicted as host query, while the rest queries as supportive queries. Supportive queries and their intermediate coarse masks, predicted from the only support image, serve as pseudo support data to facilitate segmentation of the host query. In other words, the host query is segmented by supports from one real support image and multiple pseudo ones (i.e., supportive queries), resulting in multiple predicted candidate masks. In pursuit of effective candidate mask integration, we construct an innovative MoME module, where a scene-driven router steers a flexible number of mask experts towards jointly making comprehensive decisions on candidate masks. Such module consolidates expert decisions as the final result, expanding the model capability to deal with diverse images.
@@
Note that the host and supportive queries hold an equal and interchangeable role. Each one among a group of query images can be the host query to be segmented finally, and all the query images offer favorable information mutually. Due to the need for testing many query samples in real-world scenarios, it is reasonable and feasible to collect and use various queries and apply Group-On for reciprocal help. Experiments on three standard benchmarks show that our approach significantly outperforms previous works by considerable margins, e.g., 8.2% mIoU improvement on COCO-20i [5]. Our contributions are as follows:
@@
  • We rethink the real-world application scenarios and propose Group-On, a novel and effective training approach for one-shot segmentation. For queries uploaded by users, Group-On processes category-identical ones together to address the intra-class variation issue by leveraging the rich object pattern variety.
  • We propose an innovative MoME module for candidate mask integration, where a scene-driven router guides a flexible number of mask experts to cooperate and make decisions, expanding the model capability to handle various images.
@@
  • We perform extensive experiments and analysis on three standard and challenging FSS benchmarks. The results show that the performances of one-shot segmentation baselines with our Group-On are comparable to the counterparts of conventional 5-shot segmentation.
II Methodology
@@
Figure 2: Illustrating Group-On in the one-shot setting. Different from the conventional few-shot setting, a group of query images along with one support image is input into the model. After the coarse masks of all the query images are segmented, the supportive image-mask pairs function as pseudo support to segment the host query again. Eventually, the final result is produced by the proposed MoME module, where a flexible number of mask experts make decisions on candidate masks guided by a scene-driven router. Note that the host and supportive queries play an equal and interchangeable role.
II-A Problem Setting
@@
Few-shot segmentation (FSS) aims to segment images with limited annotated examples. Following prior works [6, 7], we adopt the prevalent meta-training paradigm called episodic training [8]. Concretely, a dataset is divided into two class-disjoint sets $\mathcal{D}_{train}$ and $\mathcal{D}_{test}$, each of which contains many episodes. In an episode, there exist a support set $\mathcal{S}=\{(\textit{I}^{\text{s}},\textit{M}^{\text{s}})\}$ and a query set $\mathcal{Q}=\{(\textit{I}^{\text{q}},\textit{M}^{\text{q}})\}$ from the same class, where I and M represent raw images and their masks. Each episode instantiates an $N$-way $K$-shot segmentation task, which means selecting $K$ support image-label pairs per category from $N$ different categories to construct $\mathcal{S}$. Current approaches typically use one support image-label pair for training whereas one (i.e., 1-way 1-shot) or five (i.e., 1-way 5-shot) for testing, and one query image during both stages. The number of support images matters since it determines annotation burdens. We always seek to obtain similar performances via as few support images as possible and preferably only one. That is why many models [9, 10] freeze the number of support images to one for training.
@@
In this work, apart from the 1-way setting following [6], we propose and undertake a 1-support $N_{q}$-query setting ($N_{q}>1$) extended from the conventional settings. That is, we use ONE support image but multiple query images to help each other in the process.
II-B Group-On Process under One-Shot Setting
@@
Aiming to maximize the semantic information of different patterns in the same class as much as possible, we use a group of $N_{q}$ query images and one support image-label pair as input to the model. The pivotal competitive advantage of Group-On (see Fig. 2) is in utilizing the group of query images for mutual knowledge complement. Among the $N_{q}$ query images in the group, the final predicted image is called the host query while the rest images are supportive queries. For simple illustration, in this section we assume that $N_{q}=2$, and the host query and its feature are denoted as $\textit{I}^{\text{host}}$ and $\textit{F}^{\text{host}}$, while the supportive ones as $\textit{I}^{\text{supp}}$ and $\textit{F}^{\text{supp}}$.
@@
The entire data stream goes through the segmentation head $f$ twice. In the first time, the support feature-mask pair ($\textit{F}^{\text{s}}$, $\textit{M}^{\text{s}}$) and the query features $\textit{F}^{\text{host}},~{}\textit{F}^{\text{supp}}_{i}$’s are input into the model to generate the initial coarse masks, as:
$\textit{{C}}_{1}=f(\textit{F}^{\text{s}},~{}\textit{M}^{\text{s}},~{}\textit{F}^{\text{host}}),\ \ \ \ \ \hat{\textit{M}}^{\text{supp}}_{i}=f(\textit{F}^{\text{s}},~{}\textit{M}^{\text{s}},~{}\textit{F}^{\text{supp}}_{i}),$ (1)
@@
where $\textit{{C}}_{1}$ is a candidate mask for the host query, and $\hat{\textit{M}}^{\text{supp}}_{i}$ is the coarse mask predicted using the real support feature and its mask ($\textit{F}^{\text{s}}$, $\textit{M}^{\text{s}}$), $i=1,\ldots,N_{q}-1$ ($N_{q}=2$ here).
@@
Those predicted coarse masks $\hat{\textit{M}}^{\text{supp}}_{1}$, $\ldots$,… , $\hat{\textit{M}}^{\text{supp}}_{N_{q}-1}$ and their corresponding supportive queries are used as pseudo support data, which are input into the segmentation head again to compute the other candidate masks, as:
$\textit{{C}}_{j+1}=f(\textit{F}^{\text{supp}}_{j},~{}\hat{\textit{M}}^{\text{supp}}_{j},~{}\textit{F}^{\text{host}}),\ \ j=1,\ldots,N_{q}-1,$ (2)
@@
where $\textit{{C}}_{j+1}$ is the candidate mask of $\textit{I}^{\text{host}}$ derived from the supportive feature-mask pair ($\textit{F}^{\text{supp}}_{j},~{}\hat{\textit{M}}^{\text{supp}}_{j}$), $j=1,\ldots,N_{q}-1$. Hence, we obtain $N_{q}$ candidate masks $\textit{{C}}_{1},\ldots,\textit{{C}}_{N_{q}}$, produced from true and pseudo references in distinct appearances but from the same class, thus increasing the object diversity. The final result $\hat{\textit{M}}^{\text{host}}$ is computed by fusing these $N_{q}$ candidate masks via a novel MoME module (MoME), as:
$\hat{\textit{M}}^{\text{host}}=\texttt{MoME}(\textit{{C}}_{1},\ldots,\textit{{C}}_{N_{q}}).$ (3)
@@
It is noteworthy that under the 1-shot setting, all the query images are treated equivalently when applying Group-On. That is, the host query is not limited only to $\textit{I}^{\text{host}}$ but must be every query image, and the remaining supportive queries are processed in the same way as what is described above. This process is able to attain mutual knowledge complement. The reason for this is that it allows to make the most out of clues hidden in the query data and promote efficiency.
@@
TABLE I: Segmentation performances on the COCO-20i (left) and PASCAL-5i (right) datasets in the 1-support $N_{q}$-query setting. HSNet5 and ASNet5 represent results of the baselines HSNet (HS) and ASNet (AS) under the conventional 5-support setting, while the rest results are all in the 1-support setting. HSNet, ASNet: Our reimplementation with the CLIP backbone.
@@
5-support HSNet5 [6] 43.3 51.3 48.2 45.0 46.9 70.7 70.3 73.2 67.3 67.1 69.5 80.6
ASNet5 [1] 47.6 50.1 47.7 46.4 47.9 71.6 72.6 74.3 65.3 67.1 70.8 80.4
1-support ABCNet [11] 42.3 46.2 46.0 42.0 44.1 69.9 68.8 73.4 62.3 59.5 66.0 76.0
MCE [12] 42.1 48.3 43.7 42.76 44.2 - 65.3 71.2 66.2 61.0 65.9 78.1
HSNet [6] 36.3 43.1 38.7 38.7 39.2 68.2 64.3 70.7 60.3 60.5 64.0 76.7
ASNet [1] 41.5 44.1 42.8 40.6 42.4 68.8 68.9 71.7 61.1 62.7 66.1 77.7
Group-On (HS) 41.9 48.9 46.7 43.1 45.2 (6.0 $\uparrow$ ) 71.8 (3.6 $\uparrow$ ) 71.3 73.5 61.8 64.2 67.7 (3.7 $\uparrow$ ) 78.8 (2.1 $\uparrow$ )
Group-On (AS) 46.4 49.6 48.6 46.7 47.8 (5.4 $\uparrow$ ) 72.6 (3.8 $\uparrow$ ) 73.1 73.9 61.7 67.7 69.1 (3.0 $\uparrow$ ) 80.1 (2.4 $\uparrow$ )